# MSDet: Receptive Field Enhanced Multiscale Detection for Tiny Pulmonary Nodule

### v1

@@ Abstract
Pulmonary nodules are critical indicators for the early diagnosis of lung cancer, making their detection essential for timely treatment. However, traditional CT imaging methods suffered from cumbersome procedures, low detection rates, and poor localization accuracy. The subtle differences between pulmonary nodules and surrounding tissues in complex lung CT images, combined with repeated downsampling in feature extraction networks, often lead to missed or false detections of small nodules. Existing methods such as FPN, with its fixed feature fusion and limited receptive field, struggle to effectively overcome these issues. To address these challenges, our paper proposed three key contributions: Firstly, we proposed MSDet, a multiscale attention and receptive field network for detecting tiny pulmonary nodules. Secondly, we proposed the extended receptive domain (ERD) strategy to capture richer contextual information and reduce false positives caused by nodule occlusion. We also proposed the position channel attention mechanism (PCAM) to optimize feature learning and reduce multiscale detection errors, and designed the tiny object detection block (TODB) to enhance the detection of tiny nodules. Lastly, we conducted thorough experiments on the public LUNA16 dataset, achieving state-of-the-art performance, with an mAP improvement of 8.8% over the previous state-of-the-art method YOLOv8. These advancements significantly boosted detection accuracy and reliability, providing a more effective solution for early lung cancer diagnosis. The code will be available at https://github.com/CaiGuoHui123/MSDet.
@@ I Introduction
Lung cancer remains the leading malignancy in terms of both incidence and mortality rates worldwide [1]. Pulmonary nodules are among the earliest clinical manifestations of lung cancer, underscoring the critical importance of early detection. Currently, computed tomography (CT) is the primary modality utilized by clinicians for the screening of pulmonary nodules [2], [3]. Pulmonary nodules, typically characterized as intrapulmonary spherical lesions, range in diameter from 3 to 30 mm. The complexity of the lung environment and the variability of nodule size and morphology make nodule detection difficult. At the same time, small nodules in lung CT scans are often visually difficult to distinguish from surrounding tissues, resulting in a high rate of missed detection or false positives.
@@
Object detection methods have been widely adapted in medical imaging to automatically analyze imaging data and identify suspicious lesion areas. Two-stage detection algorithms [4, 5, 6, 7] initially detect all potential nodules from CT scans, which often results in a high number of false positives. However, these networks are typically complex and time-consuming, relying on feature maps and prior knowledge (e.g., anchor frames), leading to inaccurate localization, particularly for small or variably shaped nodules. Besides, one-stage detection algorithms [8, 9, 10, 11, 12, 13, 14] integrate nodule candidate detection and false positive reduction into a single model, achieving higher detection speeds but struggling with small nodules and occlusions, which can result in missed detections and false positives.
Meanwhile, YOLOv5, a popular one-stage detection algorithm, offers notable advantages such as adaptive anchor box computation and improved feature fusion, which enhance its speed and spatial information processing capabilities. Hence, developed a one-stage detection method based on YOLOv5 presented to be a promising trend to address these challenges. However, YOLOv5 still struggles with feature extraction, especially in complex backgrounds or with small targets. Transformers [15, 16, 17] have shown great potential in addressing these limitations. Their self-attention mechanism captures long-range dependencies and global context, making them effective for small object detection like pulmonary nodules. Incorporating attention mechanisms can thus improve feature representation and reduce both missed detections and false positives.
Inspired by these advancements, we proposed a novel one-stage detection model, MSDet, specifically designed to address the challenges of pulmonary nodule detection, including high false positive rates and low detection accuracy. Our contributions include a tiny object detection block (TODB), which captures finer feature details to improve small nodule detection, and an extended receptive domain (ERD) strategy that refines occluded nodule features to reduce false positives. Additionally, we designed the positional channel attention mechanism (PCAM) to optimize feature representation and mitigate high miss and false detection rates in multiscale nodule detection. As shown in Figure 1, MSDet achieved a significant 8.8% improvement in mAP compared to the state-of-the-art nodule detection method YOLOv8.
@@ II Related Works
Dense prediction techniques [19] have been widely adopted in medical imaging analysis [20, 21, 22, 23] particularly for tasks such as semantic segmentation [24, 25, 26, 27], and object detection [28], due to their ability to generate pixel-level predictions that enhance the precision of identifying and localizing anatomical structures or pathological regions. Detection algorithms for pulmonary nodules based on deep learning can be broadly classified into two categories: two-stage detection algorithms and one-stage detection algorithms.
@@
Two-stage detection algorithms, such as R-CNN, first generate candidate regions and then classify and localize targets. Fast R-CNN [5] accelerates detection by sharing convolutional features, while Faster R-CNN [6] improves both speed and accuracy with Region Proposal Networks (RPNs). For instance, Xu et al. [29] enhanced the Faster R-CNN model by adopting multi-scale training strategies, Online Hard Example Mining (OHEM), customized anchor sizes and ratios, deformable convolutions, and path augmentation in Feature Pyramid Networks (FPNs). This improved model significantly enhanced small target detection capabilities. Su et al. [30] proposed a Faster R-CNN-based method for lung nodule detection, achieving over 20% accuracy improvement compared to traditional algorithms, showing potential clinical value in lung disease diagnosis. Tong et al. [31] proposed a pulmonary nodule detection system using Faster R-CNN with ISODATA [32] for candidate detection and 3D-CNN with Focal Loss for false positive reduction, achieving strong results on LUNA16. However, the primary drawback of two-stage detection algorithms is the time-consuming process of generating candidate regions, which leads to slower detection speeds. Additionally, the generation of candidate regions often relies on prior knowledge, potentially resulting in inaccurate localization, especially when targets vary significantly in shape or size. These limitations hinder the efficiency and performance of two-stage detection algorithms in real-time applications.
@@
One-stage detection algorithms mitigate the limitations of two-stage algorithms by performing object detection directly on the input image without the need for candidate region generation, thus significantly accelerating the detection process. Representative one-stage detection algorithms include the YOLO series and SSD. These algorithms achieve both localization and classification in a single pass, making them highly advantageous in applications requiring real-time performance. In the context of pulmonary nodule detection, the YOLO series has been widely adopted due to its rapid detection speed and high accuracy. For example, Wu et al. [33] proposed a YOLOv7-based method for pulmonary nodule detection that incorporates Efficient Omni-Dimensional Convolution (EODConv) to enhance feature extraction precision, achieving a mAP of 94.6% on the LUNA16 dataset. Additionally, Zhanlin Ji et al. [34] improved the nodule detection performance across different scales by decoupling the feature pyramid into high-semantic and low-semantic regions to reduce semantic conflicts during fusion, achieving an accuracy of 92.3% on the Lung-PETCT-Dx dataset.
@@
However, one-stage detection algorithms face challenges due to multiple down-sampling steps, which progressively reduce the feature map size. This reduction causes the size of small pulmonary nodules on the feature map to become very small, potentially leading to the loss of fine features and making small nodules difficult to detect. Moreover, in the complex pulmonary environment, lung tissues often occlude nodules, further complicating accurate detection. These challenges result in missed detections, false positives, and high false-positive rates for occluded nodules. To address these issues, this paper proposes a novel one-stage detection network designed to integrate multiple modules specifically aimed at efficiently and accurately detecting small pulmonary nodules. This network achieves precise detection of small nodules and accurate and efficient detection of occluded and multi-scale nodules within complex pulmonary environments.
@@ III-A Overview
We proposed a novel lung nodule detection network MSDet. The overall architecture of MSDet is shown in Figure 2. Our method captures richer contextual information and reduces false positives caused by nodule occlusion through the extended receptive domain (ERD) strategy. Then, we design a dedicated position channel attention mechanism (PCAM) to effectively optimize feature learning and reduce multi-scale detection errors. Finally, the connection between different feature maps is captured by the design of the tiny object detection block (TODB), which reduces the interference of complex background tissues and aims to improve the detection accuracy of small nodules of the model. Specifically, given a lung CT image input, after convolution and feature fusion operations, multi-scale feature fusion and information transfer are realized. In the feature fusion stage, the feature map size is restored by the TODB through upsampling operation, and it is fused with feature maps of different resolutions to generate a more refined prediction feature map. Then, following the concatenation operation, the multiscale feature maps are connected and input into the ERD module designed by us, which effectively expands the receptive field of the feature map, captures more contextual information and details, and enables the network to integrate information from a larger area. As shown in Figure 3, the SPP module further enhances this capability by aggregating features at multiple scales, improving the robustness of the feature representation against variations in object size and position. A key aspect of our design is the PCAM. In the PCAM module, the position attention module captures long-range contextual information, while the channel attention module models the interdependencies between channels, thereby adjusting the attention weights according to the importance of different positions and channels. Finally, the model generates output feature maps at multiple scales to produce the final prediction results.
@@ III-B Tiny Object Detection Block (TODB)
The vanilla YOLOv5, primarily designed for handling natural images, shows limited efficacy in detecting pulmonary nodules in CT images. During network downsampling, the deepest feature map size is reduced to 20×20 pixels . In the YOLOv5, multiple downsampling operations progressively reduce the feature map size, causing small pulmonary nodules to appear extremely tiny in the feature maps. For example, a 32x downsampling reduces a 20×20 pixel nodule to 1×1 pixel, making it nearly undetectable. As a result, YOLOv5 performs poorly on pulmonary nodule detection in CT scans.
To address this challenge, we designed the TODB module to capture the relationships between different feature maps. This module aims to minimize interference from complex surrounding tissues in medical images, thereby enhancing the recognition accuracy of small nodules. By leveraging the TODB module, the network can more effectively extract and identify features of pulmonary nodules, improving detection accuracy. Figure 4 illustrates the specific structure of the TODB, which optimizes feature map connections and information extraction, significantly boosting the detection performance of small nodules as detailed below.
@@ III-C Extended Receptive Domain (ERD)
Decoupling the network’s bottleneck structures can enhance lung nodule detection but may reduce the receptive field, limiting the ability to capture global contextual information. To address this, we propose an extended receptive domain (ERD) strategy, which expands the receptive field while maintaining computational efficiency, allowing for improved detection of nodules across various sizes and scales.
As shown in Figure 6, ERD uses a multi-branch structure to capture multiscale features. It integrates three parallel 3×3 dilated convolution branches, a 1×1 convolution branch, and an identity branch. The dilated convolutions use dilation rates $R=(1,3,5)$ to provide varying receptive fields while keeping the resolution intact by setting the padding to $P=(1,3,5)$. The 1×1 convolution branch captures fine details and enhances positional information.
@@ III-D Position Channel Attention Mechanism (PCAM)
Attention mechanisms in CNNs highlight key image features while suppressing irrelevant information, improving model performance. Existing methods primarily rely on first-order statistics, limiting their ability to capture complex feature interactions. To address this, we propose PCAM, as shown in Figure 7, which incorporates higher-order statistics to improve the representation of complex activation features, and is applied to the bottleneck of the MSDet model to enhance focus on key regions in pulmonary nodule images.
PCAM consists of a positional attention module and a channel attention module. The positional attention module captures long-range spatial dependencies by encoding contextual information into local features. Given a feature map $Q\in\mathbb{R}^{C\times H\times W}$, we generate two feature maps, $R$ and $S$, and compute a spatial attention map $U\in\mathbb{R}^{N\times N}$ as follows:
@@ IV-A Dataset and Evaluation Matrices
In terms of evaluation metrics, the performance of the pulmonary nodule detection model is assessed using several standard measures. Precision quantifies the accuracy of positively predicted instances, aiming to minimize false positives. Recall evaluates the model’s ability to identify all relevant cases [35], which is crucial for thorough detection. The F1-score merges precision and recall in a single metric, reflecting the balance between detection accuracy and completeness. Average Precision (AP), calculated over the Precision-Recall curve, provides detailed insights into the model’s precision across various recall levels. Mean Average Precision (mAP) averages these AP values across different object classes and sizes, offering a global view of the model’s overall accuracy.
@@ IV-B Implementation Details
TABLE I: Ablation study of MSDet on the LUNA16 dataset, evaluating the impact of different modules (TODB, ERD, and PCAM) on detection performance. The table presents the precision, recall, and mAP0.5 results for various configurations of the MSDet model, demonstrating significant improvements when all modules are combined.
YOLOv5 TODB ERD PCAM Precision (%) Recall (%) mAP0.5 (%)
83.7 ± 0.09 81.3 ± 0.07 84.7 ± 0.03
88.4 ± 0.04 87.2 ± 0.05 89.1 ± 0.08
88.5 ± 0.02 87.6 ± 0.06 88.9 ± 0.04
87.1 ± 0.09 88.7 ± 0.05 88.3 ± 0.07
91.6 ± 0.08 90.1 ± 0.06 92.8 ± 0.02
91.4 ± 0.04 91.2 ± 0.03 92.5 ± 0.08
91.2 ± 0.09 91.4 ± 0.05 91.9 ± 0.01
97.2 ± 0.06 96.1 ± 0.05 97.3 ± 0.03
TABLE II: Detection performance comparison using different attention modules on the LUNA16 dataset. This table presents precision, recall, F1 score, and mAP results across various IoU thresholds and nodule sizes, demonstrating the superior performance of the PCAM (Ours) module compared to other state-of-the-art attention mechanisms.
The lung nodule detection workflow is depicted in Figure 8.
@@ IV-C Comparative Studies
We conducted comparative analyses using the LUNA16 lung nodule detection dataset with state-of-the-art methods, as shown in Table III. The results demonstrate that our proposed MSDet algorithm significantly outperforms various leading lung nodule detection models.
Specifically, MSDet achieved a mAP of 97.3%, surpassing several popular models. Notably, MSDet exceeded the performance of the two-stage model Faster R-CNN by 6.6 percentage points and outperformed the DLDS-CT method by 16.1 percentage points. Furthermore, compared to recent YOLO models, MSDet demonstrated improvements of 11.5%, 14.51%, 16.02%, and 8.8% over YOLOv5, YOLOv6, YOLOv7, and YOLOv8, respectively.
In addition, MSDet outperformed EfficientDet-d, CenterNet, YOLOv5-CASP, and YOLO-MSRF models, with mAP improvements of 9.9%, 12%, 25.3%, and 2.7%, respectively. These results indicate the overall superiority of MSDet in terms of detection accuracy and efficiency. The experimental findings validate the robustness and effectiveness of our proposed approach in lung nodule detection.
@@ IV-D Ablation Studies
The ablation study results in Table I demonstrate that both TODB and ERD significantly improve the performance of the Basic Model on the LUNA16 dataset. TODB alone increased precision from 83.7% to 88.4%, recall from 81.3% to 87.2%, and mAP0.5 from 84.7% to 89.1%. Similarly, ERD alone improved precision to 88.5%, recall to 87.6%, and mAP0.5 to 88.9%. When combined with other modules (ERD, TODB, and PCAM), these metrics reached their highest values: 97.2% precision, 96.1% recall, and 97.3% mAP0.5. These results highlight the significant role of TODB and ERD in enhancing detection precision, recall, and overall performance in lung nodule detection.
Figure 9: Comparison chart of mAPS–mAPL–mAP0.5. The y-axis corresponds to mAP0.5 (mean Average Precision at IoU threshold of 0.5), where higher values indicate better performance. The x-axis represents the detection performance of micro-small and relatively large pulmonary nodules. Our method, PCAM, achieved the best balance between performance and accuracy in multi-scale pulmonary nodule detection.
TABLE III: Comparison of MSDet with state-of-the-art lung nodule detection networks in terms of mAP. The table highlights the superior performance of MSDet compared to widely-used models such as YOLO, Faster R-CNN, and LungSeek, demonstrating the robustness of our approach in achieving higher detection accuracy.
@@
As shown in Table II and Table I, the PCAM (Ours) module further enhances pulmonary nodule detection. With a precision of 93.1%, recall of 92.5%, and an F1 score of 92.8%, PCAM effectively balances precision and recall while minimizing false positives. It also achieves superior mAP scores across various IoU thresholds, demonstrating robust detection performance for nodules of different sizes. The integration of position-sensitive and channel attention mechanisms enhances multi-scale detection, as illustrated in Figure 9. The ablation study confirms PCAM’s impact on the Basic Model’s performance on the LUNA16 dataset. With PCAM alone, precision increased from 83.7% to 87.1%, recall from 81.3% to 88.7%, and mAP0.5 from 84.7% to 88.3%. When combined with TODB and ERD, these metrics reached final values of 97.2% precision, 96.1% recall, and 97.3% mAP0.5, underscoring PCAM’s effectiveness in boosting detection accuracy and overall performance.

### v2

@@ Abstract
Pulmonary nodules are critical for early lung cancer diagnosis, but traditional CT imaging methods suffer from low detection rates and poor localization. Small nodule detection is challenging due to subtle differences in density and issues like occlusion. Existing methods such as FPN, with its fixed feature fusion and limited receptive field, struggle to effectively overcome these issues. To address these challenges, our paper proposed three key contributions: Firstly, we proposed MSDet, a multiscale attention and receptive field network for detecting tiny pulmonary nodules. Secondly, we proposed the extended receptive domain (ERD) strategy to capture richer contextual information and reduce false positives caused by nodule occlusion. We also proposed the position channel attention mechanism (PCAM) to optimize feature learning and reduce multiscale detection errors, and designed the tiny object detection block (TODB) to enhance the detection of tiny nodules. Experiments on the LUNA16 dataset show an 8.8% improvement in mAP over YOLOv8, achieving state-of-the-art performance. The code is available at https://github.com/CaiGuoHui123/MSDet.
@@ I Introduction
Lung cancer remains the leading cause of cancer-related incidence and mortality worldwide [1]. Pulmonary nodules, often the earliest sign of lung cancer, highlight the importance of early detection. Computed tomography (CT) is the primary method for screening pulmonary nodules. Pulmonary nodules, typically spherical lesions 3–30 mm in diameter, vary in size and morphology, making detection challenging. Small nodules are often difficult to distinguish from surrounding tissues in CT scans, leading to high rates of missed detection and false positives.
@@
Object detection methods are widely used in medical imaging to identify suspicious lesions. Two-stage detection algorithms [2] first detect all potential nodules, often resulting in numerous false positives. These methods are complex and time-consuming, relying on feature maps and prior knowledge (e.g., anchor frames), which can lead to inaccurate localization, especially for small or irregularly shaped nodules. One-stage detection algorithms [3] combine candidate detection and false positive reduction, improving speed but struggling with small nodules and occlusions, which may cause missed detections and false positives.
YOLOv5, a popular one-stage detection algorithm, enhances speed and spatial processing but struggles with feature extraction, especially for small targets in complex backgrounds. Transformers [4], with their self-attention mechanism, can capture long-range dependencies and global context, improving detection accuracy for small objects like pulmonary nodules by reducing missed detections and false positives.
Inspired by these advancements, we propose MSDet, a novel one-stage model for pulmonary nodule detection, targeting high false positive rates and low accuracy. Our contributions include a tiny object detection block (TODB) to capture finer details for small nodule detection, an extended receptive domain (ERD) strategy to reduce false positives from occlusions, and a positional channel attention mechanism (PCAM) to optimize feature representation. As shown in Figure 1, MSDet achieved a significant 8.8% improvement in mAP compared to the state-of-the-art nodule detection method YOLOv8.
@@ II Related Works
Dense prediction techniques are widely used in medical imaging, especially for semantic segmentation and object detection [6], due to their ability to make pixel-level predictions that improve precision in identifying and localizing anatomical or pathological regions. Deep learning-based pulmonary nodule detection algorithms can be broadly classified into two categories: two-stage and one-stage detection algorithms.
@@
Two-stage algorithms, like R-CNN, first generate candidate regions and then classify and localize targets. Fast R-CNN accelerates detection by sharing convolutional features, while Faster R-CNN improves both speed and accuracy with Region Proposal Networks (RPNs). Xu et al. [7] enhanced Faster R-CNN with multi-scale training, Online Hard Example Mining (OHEM), customized anchor sizes, and deformable convolutions, improving small target detection. Tong et al. [8] combined Faster R-CNN with ISODATA and 3D-CNN for false positive reduction, achieving strong results on LUNA16. However, the main drawback of two-stage methods is their slow detection speed due to the time-consuming candidate region generation, which can also lead to inaccurate localization, especially with varying target shapes and sizes.
@@
One-stage detection algorithms improve on two-stage methods by detecting objects directly from the input image, eliminating candidate region generation and accelerating the detection process. Popular algorithms like YOLO and SSD achieve both localization and classification in a single pass, making them suitable for real-time applications. In pulmonary nodule detection, the YOLO series is widely used due to its speed and accuracy. For example, Wu et al. [9] proposed a YOLOv7-based method with Efficient Omni-Dimensional Convolution (EODConv), achieving a mAP of 94.6% on LUNA16. Zhanlin Ji et al. [10] improved performance by decoupling the feature pyramid into high- and low-semantic regions, achieving 92.3% accuracy on the Lung-PETCT-Dx dataset.
@@
However, one-stage algorithms face challenges from down-sampling steps that reduce the feature map size, making small nodules harder to detect. Additionally, occlusion by lung tissues complicates detection, leading to missed detections, false positives, and high false-positive rates. To address these issues, we propose a novel one-stage detection network that integrates multiple modules for more efficient and accurate detection of small, occluded, and multi-scale nodules in complex pulmonary environments.
@@ III-A Overview
We propose MSDet, a novel lung nodule detection network (Figure 2). MSDet enhances contextual information and reduces false positives from nodule occlusion using the extended receptive domain (ERD) strategy. A position channel attention mechanism (PCAM) optimizes feature learning and reduces multi-scale detection errors. The tiny object detection block (TODB) improves small nodule detection by reducing background interference. Given a lung CT image, MSDet performs convolution and feature fusion, followed by multi-scale fusion and information transfer. The TODB restores the feature map size via upsampling, which is then fused with different resolution feature maps for refined predictions. The ERD module expands the receptive field, captures more contextual details, and integrates information from a larger area. As shown in Figure 3, the SPP module further strengthens this by aggregating features at multiple scales, improving robustness. In the PCAM module, position attention captures long-range context, while channel attention models interdependencies between channels, adjusting weights based on importance. Finally, the model generates multi-scale feature maps for the final predictions.
@@ III-B Tiny Object Detection Block (TODB)
Vanilla YOLOv5, designed for natural images, struggles with detecting pulmonary nodules in CT scans due to excessive downsampling, which reduces feature map size and makes small nodules nearly undetectable. For instance, 32x downsampling reduces a 20×20 pixel nodule to just 1×1 pixel. To address this, we introduced the TODB module, which captures relationships between feature maps and minimizes interference from surrounding tissues. By optimizing feature map connections and enhancing information extraction, TODB improves the detection accuracy of small nodules, as shown in Figure 4.
@@ III-C Extended Receptive Domain (ERD)
Decoupling network bottleneck structures can improve lung nodule detection but may reduce the receptive field, limiting global context. To address this, we propose an extended receptive domain (ERD) strategy, which expands the receptive field while maintaining efficiency, enhancing detection across various nodule sizes. As shown in Figure 6, ERD uses a multi-branch structure with three parallel 3×3 dilated convolution branches (dilation rates $R=(1,3,5)$), a 1×1 convolution branch, and an identity branch. This approach provides varying receptive fields, captures fine details, and preserves positional information.
@@ III-D Position Channel Attention Mechanism (PCAM)
Attention mechanisms in CNNs highlight key features while suppressing irrelevant ones, improving performance. Existing methods rely on first-order statistics, limiting feature interaction capture. To address this, we propose PCAM (Figure 7), which incorporates higher-order statistics for better feature representation. Applied at the MSDet bottleneck, PCAM enhances focus on key regions in pulmonary nodule images. It consists of a positional attention module, which captures long-range spatial dependencies by encoding contextual information. Given a feature map $Q\in\mathbb{R}^{C\times H\times W}$, we generate two feature maps, $R$ and $S$, and compute a spatial attention map $U\in\mathbb{R}^{N\times N}$ as follows:
@@ IV-A Dataset and Evaluation Matrices
The experimental data were obtained from the publicly available LUNA16 lung CT dataset, with detailed dataset configurations provided in Supplementary Section 1. The model performance was evaluated using standard metrics such as precision, recall, F1-score, and mAP.
@@ IV-B Implementation Details
The MSDet model was implemented using PyTorch 1.13.1 and trained on an NVIDIA GPU with a batch size of 8 for 200 epochs. Detailed experimental parameters and configurations are provided in Supplementary Section 2. The detection workflow is shown in Figure 8.
@@ IV-C Comparative Studies
We performed comparative analyses on the LUNA16 dataset, as shown in Table I, against state-of-the-art methods. MSDet significantly outperforms leading models, achieving a mAP of 97.3%, surpassing Faster R-CNN by 6.6 percentage points and YOLOv5 by 11.5%. Additionally, MSDet outperformed EfficientDet-d and CenterNet by 9.9% and 12%, respectively. These results highlight MSDet’s superior detection accuracy and efficiency, validating its effectiveness for lung nodule detection.
@@ IV-D Ablation Studies
The ablation study in Supplementary Table (1) shows the significant improvements achieved by TODB, ERD, and PCAM in the model’s performance on the LUNA16 dataset. The combination of these modules led to a substantial boost in precision, recall, and mAP0.5, with final metrics reaching 97.2%, 96.1%, and 97.3%, respectively. Supplementary Table (2) compares various attention mechanisms, revealing that PCAM outperforms others in precision, recall, and F1 score, while enhancing multi-scale detection and reducing false positives. For detailed results, please refer to Supplementary, Section 3.
TABLE I: Comparison of MSDet with state-of-the-art lung nodule detection networks in terms of mAP. The table highlights the superior performance of MSDet compared to widely-used models such as YOLO, Faster R-CNN, and LungSeek.